OpenStudy (highschoolmom2010):

Verify the identity. (1 + tan^2u)(1 - sin^2u) = 1

hartnn (hartnn):

hi @highschoolmom2010, what have you tried, there are several ways using trigonometric identities in which you can prove this. One of the easier ways, i can think of is using: \(1+\tan^2x = \sec^2 x \\ \sin^2x +\cos^2 x = 1 \\ \sec x \cos x = 1\) with these 3 identities, that will be easily proved.

Directrix (directrix):

You will get a workout on trig identies with this problem: (1 + tan^2u) = sec ² u (1 - sin^2u) = cos ² u sec ² u and cos ² u are reciprocal trig functions Putting this together: (1 + tan^2u) (1 - sin^2u) = 1 sec ² u * cos ² u = 1 See if this makes sense: @highschoolmom2010 1 = 1 @